OpenStudy (anonymous):

One card is selected at random from a deck of cards. Determine the probability that the card is a 5 or a 7

OpenStudy (anonymous):

A common deck has 52 cards, containg totally 4 5s and 4 7s. So Chance = 4+4 / 52 => 8/52 => 2/13 or 15,38 %

OpenStudy (anonymous):

There is 52 cards in a deck .. There are 4 of each type of card

OpenStudy (anonymous):

P(getting a 5) = 4/52 P(getting a 7) = 4/52 So, probability of both 5 or 7: \[P = \frac{4}{52} + \frac{4}{52} = \frac{8}{52} = \frac{2}{13}\]

OpenStudy (anonymous):

All of you have been very helpful....not so good at math and its very hard to find help...thanks for everything

OpenStudy (anonymous):

Welcome dear..